### v2

# Adaptive Collision Sensitivity for Efficient and Safe Human-Robot Collaboration

Abstract
What is considered safe for a robot operator during physical human-robot collaboration (HRC) is specified in corresponding HRC standards (e.g., ISO/TS 15066). The regime that allows collisions between the moving robot and the operator, called Power and Force Limiting (PFL), restricts the permissible contact forces. Using the same fixed contact thresholds on the entire robot surface results in significant and unnecessary productivity losses, as the robot needs to stop even when impact forces are within limits. Here we present a framework that decides whether the robot should interrupt or continue its motion based on estimated collision force computed individually for different parts of the robot body and dynamically on the fly, based on the effective mass of each robot link and the link velocity. We performed experiments on simulated and real 6-axis collaborative robot arm (UR10e) with sensitive skin (AIRSKIN) for collision detection and isolation. To demonstrate the generality of our method, we added experiments on the simulated KUKA LBR iiwa robot, where collision detection and isolation draws on joint torque sensing. On a mock pick-and-place scenario with both transient and quasi-static collisions, we demonstrate how sensitivity to collisions influences the task performance and number of stops. We show an increase in productivity over 45% from using the standard approach that interrupts the tasks during every collision. While reducing the cycle time and the number of interruptions, our framework also ensures the safety of human operators. The method is applicable to any robot for which the effective mass can be calculated.
I Introduction
Robots are leaving safety fences and start to share workspace or even living space with humans. As they leave controlled environments and enter domains that are much less structured, they need to dynamically adapt to unpredictable interactions and guarantee safety at every moment. There has been rapid development in this field of physical human-robot interaction (pHRI) or human-robot collaboration (HRC) in the last decade (see [1, 2, 3] for surveys), a growing market of collaborative robots, along with the introduction and revision of new safety standards ([4, 5, 6]). According to [6], there are two main ways of satisfying the safety requirements when a human physically collaborates with a robot: (i) Power and Force Limiting (PFL)
and (ii) Speed and Separation Monitoring (SSM). For PFL, physical contacts with a moving robot are allowed, but the forces, pressures, and energy absorbed during a collision need to be within part-specific limits of the human body. This translates into lightweight structure, soft padding, and no pinch points on the robot side, in combination with robot interaction control methods for the post-impact phase: collision detection and response relying on motor load measurements, force/torque, or joint torque sensing (e.g., [8]). However, the speed and payload of robots that comply with this safety requirement is limited.
Furthermore, in practice, exploiting the full potential of the PFL regime, i.e., stopping the robot only if the prescribed force threshold is exceeded, is difficult or impossible. According to [7], a collision needs to be 1) detected, 2) isolated (which robot link)), 3) identified (estimating the external torque or force), 4) classified, and 5) then a robot reaction is triggered—see Fig. 2. The accuracy or even the possibility of performing some of these steps depends on the type of contact (single or multiple), the contact location, and in particular on the sensory equipment of the robot. Electronic skins covering large parts of the robot surface [9, 10, 11, 12] facilitate detection and isolation. For identification, force/torque or joint torque sensors together with accurate models of the robot dynamics are the best. However, correct estimation of forces, pressures, and energies on the body of the human operator is not possible for several reasons. First, only the first very brief evolution of the collision is used for collision identification. According to PFL, the first should be considered for transient contact and even longer for quasi-static contact. Second, even if external force on the robot body can be estimated, applying Newton’s third law and assuming that the same force acts on the human body is problematic. Third, without additional visual sensors, the robot does not know with which part of the body it is colliding (different limits apply and different tissue stiffness is needed to model the collision). In practice, collaborative robot applications are thus configured too conservatively, i.e., stopping the robot as soon as a collision is detected. Frequent stops of the robot then negatively impact productivity.
In this work, we present a new method that aims to maximize productivity in the PFL regime as illustrated in Fig. 2. The idea is to perform collision detection and isolation using a robotic skin (or joint-torque sensors), “identify” the force using a simple equation and trigger a robot reaction. We tested the method on two robots, and the collision is detected and isolated using a pressure-sensitive electronic skin (AIRSKIN) covering the entire robot body of the UR10e robot (both real and simulated), or using joint-torque sensors of the KUKA LBR iiwa 7 robotic arm (only simulated). The key innovation here is that we online estimate the collision force from the robot velocity and effective mass and then use that estimate to decide how to react to the collision. If the estimated force is low, the robot will react by continuing the execution of the task. If the estimated impact force is high, the robot will stop.
To our knowledge, this is the first study that adaptively sets collision thresholds on a collaborative robot that brings safety (force limits according to [6]) and productivity under one hood. Solutions for use in industry such as AIRSKIN normally employ the same threshold in all skin pads and use a highly sensitive setting, resulting in too frequent task interruptions. In research articles employing large-area skins, same thresholds are also typically employed (e.g., [13]).
Our recent work on this topic [14] introduced this idea, but was largely based on simulation, featured only preliminary real robot experiments and was not safety focused. Here, we present systematic experiments on a real robot by utilizing a robust environment allowing for repeatable measurements. In addition, the collision handling has been redesigned to be safer. Previously, the collision was detected only if the pressure applied to the robot was higher than a predefined threshold. Now, every collision is detected, and only the reaction is based on the estimated force. The new experimental evaluation contains not only transient but also quasi-static (clamping) contacts. We added a second robot with different dimensions and physical properties, and different collision-sensing technology. Finally, all experiments were performed with multiple Cartesian velocities.
II Method
The core of this work is the evaluation of simulated and real-world impacts in a setup that simulated human-robot cooperation with several obstacles during a task performed by the robot. This section describes the theoretical background, setup, and robots used in the experiments.
II-A Task and Environment
The environment (Fig. 3) we prepared to carry out the experiments consists of three buckets (B1, B2, B3) hung on a rope from the ceiling. The buckets are filled with small stones with a total weight of . The weight was selected according to the average weight of the whole human arm from ISO/TS 15066[6]. The buckets can freely move in space, thus simulating the impact of the robot with the human arm. The impacts with buckets are transient, that is, dynamic impacts in which a human body part is impacted by a moving robot and the human can recoil or retract [6]. The setup also includes the CBSF-75-Basic certified measuring device for validation of collaborative tasks with the range of . This device is firmly attached to the table and is used to simulate a clamping scenario. This impact is quasi-static, that is, impact of a human body part with the robot, when the robot clamps or crushes the human for a prolonged period of time [6].
To be able to evaluate the effect of different collision threshold settings on performance, we define a mock pick-and-place task—see the video at https://youtu.be/UyOa--6Jpj4. The direction of the movements is depicted in Fig. 3. The order of movements starts from the top green line (movement in the y-axis of the robot) and collision with the first bucket B1, continues with slight downward movement (negative z-axis) and movement in the robot’s x-axis. Then it continues with impact to B2. The robot then goes again downward to collide with the measuring device, and returns up after the collision. Finally, the robot moves in its negative y-axis and collides with bucket B3.
II-B Robot, Skin and Software
We have used two robots.
Universal Robots UR10e manipulator with Airskin. This robot is a 6- Degree of Freedom (DoF) robotic manipulator and the arm itself is designed for collaborative purposes. Our robot is further equipped with robotic skin AIRSKIN111https://www.airskin.io/. Using a special Airskin maintenance controller, the pressure values from individual skin parts (also called pads; 11 of them placed on our robot) can be retrieved at approximately . Different from standard Airskin deployment in the industry where any detected collision stops the robot, here we use the pressure sensor for collision detection and isolation (which robot link collided) and choose to stop or continue using our algorithm.
A real and simulated version of the robot was used.
KUKA LBR iiwa 7. The second robot used is also collaborative. It is equipped with joint torque sensors in each of its joints and can also estimate the external torques, allowing collision detection.
A simulated version of the robot was used.
Both robots are integrated and controlled using the Robot Operating System (ROS) middleware. The task performed by the robots (see Section II-A) is defined in the Cartesian space, but we utilize control in the joint velocity space as it gives us more control over the movements and contact reactions.
Simulation. We used simulators that we developed (available at https://rustlluk.github.io/adaptive-sensitivity/). For both robots, the underlying simulation of physics is achieved with PyBullet [15]. The simulator also contains the buckets and force-measuring device with realistic physical behavior allowing us to do a higher number of experiments than in the real world.
II-C Collision Handling Strategies
We employed three different strategies to handle collision events: (i) FACTORY; (ii) FIXED MASS; (iii) ADAPTIVE MASS.
II-C1 FACTORY
The first option is common (“industrial”) strategy used with collaborative robots in industry: Whenever a contact is detected, the robot is stopped immediately. This serves as a baseline to gauge how much productivity can be gained when a better collision reaction strategy is used.
Dynamic collision handling strategies. The two strategies described below (FIXED MASS and ADAPTIVE MASS) are dynamic, that is, the robot reacts to a collision by stopping or continuing based on properties computed on the fly.
Both methods are based on equation from ISO/TS 15066 [6], which defines the maximal allowed robot velocity given the permissible collision force limit as
where is the spring constant of impacted human body, is the Effective Mass (EF) of the robot, is the EF of the impacted human body, is the Cartesian velocity of the robot link and is the maximal permissible force.
For this work, we need to rearrange the equation and express
We evaluate an instance of this equation for every robot link on the run, plugging in the instantaneous velocity of every robot link. If collision of a particular robot link is detected, the equation is used to estimate the impact force.
The spring constant is fixed in our experiments to corresponding to the spring constant of the back of the non-dominant hand as suggested by ISO/TS 15066. We selected the mass as (and we made the hanging buckets weigh the same) based on the average weight of human arm defined also in [6].
The two methods then differ in how they determine the robot mass .
II-C2 FIXED MASS
The mass is in ISO/TS 15066 defined as Half of Moving Mass (HMM) of the robot. In our experiments, this weight is computed as half of the sum of masses of links up to the link in collision.
II-C3 ADAPTIVE MASS
This policy also uses Eq. 2 to compute the permissible force. However, now is actually the real Effective Mass (EF) of the robot. EF is a property of the robot (each point of the robot), which can be computed using the dynamics of the robot [16]. EF is always computed in a given direction. We consider this direction to be the direction of the link on which the collision is detected. Using this direction , one can compute the EF as
where is the upper matrix of Cartesian kinetic energy matrix , that can be computed as
where is the joint space inertia matrix.
This property of robots is well known in the community, but is arguably still not used often. The reason is its sensitivity to proper dynamic model of the robot and instability in some configurations (mainly near the base of the robot; see [17] for discussion on the topic). We provide some examples in Fig. 4, where the EF in every configuration of our task is shown for three skin pads (the pads that can be in collision with any of the obstacles). We can see that the EF is in most cases lower than HMM, which was the motivation to use this property in our dynamic threshold setting policy. However, we can also see that it is higher in the beginning of the task for one of the pads. The given skin part (grey pad in Fig. 1) is in the beginning of the movement very close to the robot base and is moving in such a way that it realistically cannot exert any force to obstacles (see the video at https://youtu.be/UyOa--6Jpj4).
Setting collision force thresholds. For both of the dynamic collision handling strategies, on collision of a particular robot link, the collision force is estimated based on Eq. 2 and compared to force thresholds obtained from [6]: and for quasi-static and transient contacts, respectively.
The collision reaction is thus defined as:
for transient contacts and
for quasi-static contacts.
II-D Transient Contact Force Estimation
In physical HRI, safety is at stake and we need to make sure that the dynamic estimates of the impact forces, which are used to boost productivity here, are accurate or conservative (actual collision forces are always lower than estimated). In our setup, the certified force-measuring device is used only at the location of the quasi-static contact. In order to empirically evaluate also the forces of the transient collisions with the buckets suspended from the ceiling, we measured their deflection after impact and approximated them as pendulum. Therefore, we added a large protractor (circa A4 paper) behind the first bucket B1 and captured the swing of the bucket using a camera and extracted the maximal angle of the pendulum . We know that potential energy and kinetic energy , where is the weight of the bucket, is the length of the string, is the angular velocity and height . Using the law of conservation of energy, the energies at the maximal angle should be equal, i.e., , and from there we can solve for as
For this calculation, we used the baseline, FACTORY, setting where the robot stops immediately after impact. We assume that all momentum will be transferred to the bucket, i.e., , with momentum of the bucket given by . From impulse-momentum principle we know that the average force over a small period of time is equal to the moment, i.e., . If we put all together and solve for we get
III Experiments and Results
We performed several simulated and real-world experiments using our setup that mimics possible collision scenarios in human-robot collaboration. All data collected during the experiments are available at https://rustlluk.github.io/adaptive-sensitivity/.
III-A Experiments Settings
As described in Section II-A, the robot performed a mock pick-and-place task in a controlled environment with three buckets ( each) suspended from the ceiling that acted as obstacles (simulation of a collision with the human arm). Collisions were detected and isolated and depending on the collision reaction strategy, the robot was commanded to stop or to continue task operation. If stop was triggered, the robot remained stopped for . If there was no longer a collision after the waiting period, the robot continued its task. If a new collision occurred at the same place (usually the bucket swinged back to the robot), the robot evaluted it, possibly stopping again. In addition to the buckets, there was an impact measuring device attached to the table, which simulated a clamping scenario between the robot and a human operator. As this contact was not transient, if a stop was triggered, the robot stopped for and then was commanded to continue the task.
The robots were controlled in joint velocity so that the end-effector moved with a predefined Cartesian velocity. However, we had to make the movement towards the quasi-static collision at max in order not to trigger the internal safety controller of the robot.
In simulation, we tested both robots with Cartesian velocity of 200, 400, and . For each velocity and each collision strategy (FACTORY, FIXED MASS, ADAPTIVE MASS) we ran 100 repetitions of the experiments. In the real world, we tested the UR10e robot with velocities of and and performed 10 repetitions for each velocity and collision reaction strategy.
III-B UR10e with AIRSKIN
We will start with the UR10e robot, for which we performed experiments in both the simulation (left side in Fig. 5) and the real world (right side in Fig. 5). The first, and arguably most important, results in both worlds is the difference in task completion time between different collision handling strategies for the Cartesian velocity of . As expected, the FACTORY mode, where every touch results in a stop of the task execution, is the slowest. It is followed by the FIXED MASS and ADAPTIVE MASS modes. In the real world, the difference between the FACTORY and ADAPTIVE MASS modes is, on average, more than 6 s (45% faster; 14.5 s vs. 8 s) while FIXED MASS is closer to ADAPTIVE MASS with 10 s. In the simulation, the times are closer to each other with 11.5, 9, and 7.5 s, respectively. The discrepancy is probably cased by the physics of the buckets in the simulation. We tried to set the parameters as closely to the real world as possible (see the video at https://youtu.be/UyOa--6Jpj4), but the behavior is still not identical.
The completion time of the task is correlated with the number of stops of the robot during the task, shown in the bottom part of the Fig. 5. We can see that the trend in differences between the individual modes is the same as for the task time. In the real world, the robot stopped using ADAPTIVE MASS only half of the times (on average) compared to the FACTORY mode (4 vs. 8 stops), while FIXED MASS is again between them with 6 stops. This behavior is supported by Fig. 1, where, mainly for bucket B2, the ADAPTIVE MASS strategy is more permissive. The average number of stops is lower in the simulation with 4.5, 3.5, and 2.5 for FACTORY, FIXED MASS and ADAPTIVE MASS, respectively.
With lower velocities, there is basically no difference between ADAPTIVE MASS and FIXED MASS, as the velocities are too slow to exceed the stop thresholds for either of the strategies. We can see that for the velocity of the robot stopped on average only once during the task for both modes and both worlds. And for the velocity of it actually never needed to stop. In the FACTORY mode, the task time (and number of stops) is high. For example, in the real world with the velocity of the difference in task time is 21 s vs 6.3 s.
The results show that using ADAPTIVE MASS or FIXED MASS makes task completion faster with fewer interruptions. There is thus a clear increase in productivity. We need to verify that this does not compromise safety. In other words, we need to verify that the impact force estimates are conservative. Therefore, we utilized the CBSF-75-Basic certified measuring device and measured the forces excerted by the robot during quasi-static contact (the limit based on ISO/TS 15066 is 140 N). The results are shown in Fig. 6. The histogram is created from all 60 measurements in the real world and all were measured with a velocity of . We can see that more than 90% of the forces is lower than the force estimated using ADAPTIVE MASS (red line; 175 N). The measuring device is sensitive to the “accuracy” of the impact (it should be in the middle and in the direction of the spring of the device) and thus the variance of the measurements is high (from 40 N to 215 N), but the majority of forces are under 160 N. In a real deployment scenario, such empirical measurements would be used to adapt the dynamic thresholds, making sure the prescribed limits are always respected.
As described in Section II-D we also attempted to measure the impact force exerted during transient contacts, using a pendulum approximation. We extracted the maximum angles after collision from the real videos of a protractor behind the bucket B1. We used only values from experiments with FACTORY strategy, as it always stopped and it was easier to extract the angles. The measured angles were and for the velocities of and , respectively. The forces estimated using ADAPTIVE MASS near bucket B1 for the velocities are 190 and 280 N, respectively. Pendulum forces computed using Eq. 8 for different contact durations are shown in Table I. It is complicated to estimate the correct , but from the recordings we know that it is more than one frame and less than two, that is, 40-80 ms. We can see in Table I that the forces for 40 and 50 ms are low. However, we must take into account that these are average forces exerted throughout the duration of contact. There will probably be a higher initial force peak exerted for a shorter time duration. As both the bucket and AIRSKIN are slightly deformable, a contact duration of around 20-30 ms is a reasonable estimate. As we can see, the forces are still lower than the forces calculated using ADAPTIVE MASS.
| Contact Duration [ms] | ||||
| Velocity | 20 | 30 | 40 | 50 |
| 400 | ||||
| 600 | ||||
Based on the fact that both quasi-static and transient estimates are lower than the actual forces, we can say that ADAPTIVE MASS is effective and also still conservative, and thus safe.
III-C KUKA LBR iiwa 7
To demonstrate the general applicability of our method—it can work on any robot capable on collision detection and isolation, we also evaluated the task on a simulated KUKA LBR iiwa 7 robot. This robots does not have sensitive skin, but each of its joints is equipped with joint torque sensors that allow to detect and isolate collisions. We ran an adapted version of the same task with buckets at different places as the robot is smaller than the UR10e.
The results are shown in Fig. 7. We can see that the trends are similar to those for the UR10e. At the highest velocity () the task time is the highest for FACTORY mode (11 s) followed by FIXED MASS (10 s) and ADAPTIVE MASS (8 s), resulting in a speedup of about 30% when using ADAPTIVE MASS. The number of stops again corresponds to the task time with 4.5, 4, and 3 touches, respectively. Interestingly, at lower velocities the ADAPTIVE MASS is slower (and stops more) than FIXED MASS. It shows that FIXED MASS can actually underestimate the situation in some cases and the force computed using ADAPTIVE MASS is higher, which may be an important aspect of safety.
IV Conclusion, Discussion, and Future Work
We introduced three different collision handling strategies: One static and commonly used in the industry, where every detected collision triggers a stop of the robot, and two dynamic strategies that are calculating estimated contact force on the fly to choose how to react to the collisions—stop or continue operation. We tested the strategies using two robots in simulation (UR10e with artificial skin AIRSKIN and KUKA LBR iiwa 7 with joint-torque sensors) and with UR10e robot in the real world. The results show that when utilizing one of our dynamic methods (ADAPTIVE MASS), which calculates the estimated force using a standardized equation from ISO/TS 15066 [6] and Effective Mass (EF) of the robot, the task time can be significantly lower (up to 70% with the velocity of or 45% with , for the UR10e robot). The strategies also result in different number of interruptions during the task, which can be crucial for the health of the robot and its components.
The results are consistent for the two different robots we used, each of them using a different collision detection and isolation strategy (sensitive skin vs. joint torque sensing). Other collision detection methods are used by collaborative robots (a force/torque sensor in the robot base or using motor currents). Our method is potentially applicable to any of them, provided that collision detection and isolation (which robot link collided) is available. For the strategy that increases productivity the most (ADAPTIVE MASS), a code must be developed to estimate the effective mass of every robot link online.
We compared the collision forces estimated by our method with empirical measurements to make sure that the increase in productivity does not compromise safety. Some empirical measurements will always be indispensable before real deployment (see [18, 12] for extensive empirical measurements on the two robots used in this work).
In this work, only the simplest reaction to a collision that exceeds a pressure threshold was applied: stop the robot. Post-collision reactions can have several forms (see, for example, [19]) and may also depend on what information is available about the collision (see collision identification in [7]) and on the robot capabilities (e.g., whether torque control is available [8]). In the setup used here—with coarse isolation of the collision location in space, a pressure value reading, and a velocity control mode—one possibility is to implement an avoidance reflex that will move the colliding robot link in the opposite direction (see [14]). Artificial skins with higher spatial resolution make it possible to move away the collision point along the normal to the robot skin at that location [20].
Our method is usable for contacts with parts of the robot that are not sharp or dangerous otherwise. However, the forces and possible injuries change based on the colliding part of the robot (see [21]), which should be incorporated into the pipeline.